# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - AutoML from Basics to State-of-the-Art Techniques

### v2

@@ 7.2.1 Information Gain
Information gain is a key concept in decision tree learning, where it is used to select the attribute that best separates the data into distinct classes [57]. It is based on the concept of entropy from information theory, which measures the level of disorder or impurity in a dataset. A split that results in the greatest reduction in entropy is considered the best, as it leads to a more organized and homogenous distribution of classes [61].
@@
The term entropy has its roots in thermodynamics, where it was used to describe the amount of disorder or randomness in a physical system. The concept was later adapted by Claude Shannon in the 1940s to lay the foundation for information theory, which deals with the transmission, compression, and processing of data [62]. Shannon defined entropy as a measure of uncertainty or impurity in a system of information.
In the context of machine learning, entropy quantifies the uncertainty in predicting the class label of an instance in a dataset. If a dataset is perfectly homogeneous (i.e., all instances belong to the same class), the entropy is zero, indicating no uncertainty. On the other hand, if the dataset is evenly split between two or more classes, the entropy is at its maximum, indicating a high degree of uncertainty in classification [63].
@@
$Information\ Gain=Entropy(Parent)-\sum\left(\frac{|Child|}{|Parent|}\right)\times Entropy(Child)$
@@
In Shannon’s framework, the goal was to quantify the amount of uncertainty in a message. If you are transmitting a message and the outcome is highly predictable, then little information is gained from receiving it. However, if the outcome is highly uncertain, then the message carries more information. Shannon defined entropy mathematically to quantify this uncertainty [62]. In this way, the more uncertain or unpredictable a message is, the more ”information” it contains [64].
@@ 7.2.2 Gini Index
The Gini index is a measure used in decision trees, particularly for classification tasks, to evaluate the quality of a split. It assesses the degree of impurity or homogeneity of a node, indicating how well the split separates the data into distinct classes [58]. The Gini index is calculated as:
@@
  • $A$ is the area between the line of equality and the Lorenz curve.
@@ 7.3 Working Principles of Random Forests
Random Forests are an ensemble learning method that improves decision trees by constructing multiple trees during training and outputting the average prediction of all trees for regression, or the majority vote for classification [44]. By using multiple trees, the random forest reduces the risk of overfitting and increases the model’s generalization ability [65].
@@ 8.1 Overview of Boosting Algorithms
Boosting is a powerful ensemble learning technique that combines the predictions of several weak learners to create a strong model [66]. Unlike bagging (such as random forests), where individual models are trained independently, boosting builds models sequentially. Each subsequent model attempts to correct the errors made by the previous ones [67]. The goal is to reduce bias, making boosting an effective method for improving model accuracy, especially with complex datasets [68].
Boosting works by assigning higher weights to misclassified examples, forcing subsequent learners to focus more on these difficult cases. Boosting algorithms have been widely used in competitions like Kaggle due to their high performance, and they are known for being robust to overfitting when tuned correctly [45].
@@ 8.2 XGBoost
XGBoost stands for eXtreme Gradient Boosting. It is one of the most efficient and scalable implementations of gradient boosting algorithms [45]. XGBoost has become highly popular due to its speed, performance, and the extensive control it gives over the boosting process.
@@ 8.3 LightGBM
LightGBM (Light Gradient Boosting Machine) is another boosting framework that is optimized for efficiency and speed. It was developed to handle large datasets quickly with lower memory usage [69]. LightGBM is known for its ability to handle categorical features natively and for its leaf-wise tree growth strategy, which often results in better accuracy.
@@ 8.4 CatBoost
CatBoost [70] is a boosting algorithm developed by Yandex, and it is specifically designed to handle categorical data efficiently. Unlike XGBoost and LightGBM, CatBoost doesn’t require one-hot encoding for categorical features, making it highly effective when working with categorical data.
@@ 9.1.2 Examples of Sparse Models
  • Lasso Regression: A type of linear regression that uses an L1 penalty to shrink some coefficients to zero, effectively removing those features [49].
  • Elastic Net: Combines L1 and L2 regularization, creating a balance between ridge regression and lasso regression [71].
@@ 9.2 Principles of Group Lasso
Group Lasso [72] is a regularization technique that extends the concept of Lasso by considering groups of features instead of individual ones. It applies L1/L2 norms to groups of variables, leading to entire groups being selected or removed from the model. This is especially useful when the features naturally form groups, such as polynomial features or features derived from categorical variables.
@@ 10.1 Concept of RiskSLIM
RiskSLIM (Risk-Supersparse Linear Integer Models) is a machine learning model designed for tasks that require risk minimization, particularly in contexts where interpretability is critical [73]. Unlike many complex machine learning models that focus on maximizing prediction accuracy at the expense of transparency, RiskSLIM emphasizes both simplicity and predictive performance. It produces highly interpretable scoring systems, often represented as linear models with integer-valued coefficients. These models are particularly useful in domains like healthcare, finance, or law, where decision-making is sensitive, and understanding the model’s reasoning is as important as the accuracy itself [74].
The core idea of RiskSLIM is to balance risk minimization with the constraints of producing sparse, simple models. This balance is achieved by optimizing a loss function while enforcing constraints on the coefficients of the model. The coefficients are typically constrained to be small integers, making the final model easy to interpret and apply, even by non-experts [75].
RiskSLIM is ideal when you need to make decisions based on a risk score. For instance, in a medical setting, doctors might use a RiskSLIM model to assess the likelihood of a patient developing a condition based on various health indicators. The output would be a simple, interpretable risk score that directly correlates with the probability of the condition [76].
@@ 11.1 Basics of GridSearchCV
Grid search is one of the most popular techniques for hyperparameter tuning in machine learning [77]. It involves exhaustive searching through a manually specified subset of the hyperparameter space of a learning algorithm. When using GridSearchCV, we systematically work through different combinations of parameter values, cross-validating as we go to determine which combination gives the best performance [22].
This method is particularly useful when the search space is small and manageable, as it guarantees that all possible combinations are tested. However, for very large search spaces, more efficient techniques like random search [78] or Bayesian optimization [79] may be preferred.
@@ 12.1 Concept of AutoML
Automated Machine Learning, commonly referred to as AutoML, is the process of automating the end-to-end process of applying machine learning (ML) to real-world problems. Traditionally, applying machine learning techniques [80] to a problem requires expertise in data science, programming, and the ability to design machine learning models. However, AutoML aims to simplify this process, making it more accessible to non-experts and reducing the time required to build ML models [3].
@@
14class SimpleNet(nn.Module):
@@
17 self.fc1 = nn.Linear(10, 50)
18 self.fc2 = nn.Linear(50, 2)
@@ 12.2.1 Milestones in the Development of AutoML
  • Early 2010s: The first generation of AutoML tools emerged. Tools such as Auto-WEKA (2013) and Auto-sklearn (2015) [4] provided early platforms for automating the machine learning pipeline. These focused on automating the model selection and hyperparameter tuning processes.
  • 2018 - Neural Architecture Search (NAS): AutoML evolved beyond classical ML into deep learning with the introduction of NAS. Google introduced NASNet, which automated the process of searching for the best neural network architecture [81]. This was a significant breakthrough in designing efficient neural networks.
  • 2020 - Democratization of AutoML: Many cloud providers began offering AutoML as a service. Tools such as Google AutoML, Azure AutoML, and Amazon SageMaker Autopilot simplified the process of applying AutoML on scalable cloud infrastructure [82].
  • Present: Recent advancements in AutoML focus on efficiency, model explainability, and lowering the compute cost of model selection. Research has also been growing around topics like Few-Shot Learning, Zero-Shot Learning, and Hyperparameter Optimization [10].
@@ 13.1 Introduction to TPOT
TPOT (Tree-based Pipeline Optimization Tool) is an open-source library designed to automate the process of machine learning, with a particular emphasis on automating feature engineering, model selection, and hyperparameter tuning [83]. It utilizes genetic programming to search for the best possible machine learning pipeline by evaluating different models and combinations of preprocessing steps, features, and hyperparameters.
Machine learning can often be challenging for beginners, as it requires not only a good understanding of the data but also selecting the right models and hyperparameters. TPOT simplifies this process by automatically searching through a range of models and pipeline combinations, thus allowing users to focus more on understanding their data and less on the technical details of model selection and tuning [84].
TPOT is built on top of the popular machine learning library scikit-learn [22], and it integrates seamlessly with PyTorch for deep learning tasks. TPOT handles the process of pipeline optimization by evolving the best possible model pipeline over a series of iterations. Each iteration involves generating new pipelines, evaluating their performance, and selecting the best candidates for further optimization. This process continues until a pre-defined number of generations is reached or the optimal solution is found [85].
@@ 14.1 Introduction to AutoGluon
AutoGluon [86] is an open-source toolkit designed to simplify machine learning model development. It automates the process of building, tuning, and deploying models by handling most of the complex steps automatically. This makes it an excellent choice for beginners and professionals alike, as it reduces the need for manual tuning of hyperparameters or selecting the best model. Instead, AutoGluon focuses on performance and ease of use, allowing developers to achieve state-of-the-art results without extensive knowledge of machine learning.
@@ 15.1 H2O AutoML
H2O AutoML [87] is an open-source machine learning platform that automates the entire model training process. It’s known for providing an easy-to-use interface for training and tuning models without needing in-depth expertise in machine learning. With H2O AutoML, you can automatically train a large variety of models, including deep learning, tree-based models, and ensembles. The tool is built to handle both regression and classification problems.
@@ 15.2 MLBox
MLBox is a Python-based AutoML library designed to automate the end-to-end process of machine learning, from preprocessing to model training and optimization [88]. It emphasizes simplicity and automation, making it a great choice for beginners. It also provides powerful preprocessing capabilities, especially for handling missing data and unbalanced datasets [89].
@@ 15.3 Auto-sklearn
  • Built on scikit-learn: Uses the familiar scikit-learn API [22], making it easy to integrate into existing Python workflows.
@@ 15.4 FLAML
FLAML (Fast and Lightweight AutoML) [90] is a lightweight AutoML library developed by Microsoft, focusing on fast and efficient hyperparameter optimization. FLAML is designed to be computationally efficient, making it suitable for users who want to train models quickly without consuming significant computational resources. It is a great tool for handling both classification and regression tasks with a focus on performance and speed.
@@ 16.1 Introduction to DataRobot
DataRobot [91] is a powerful cloud-based AutoML (Automated Machine Learning) platform that allows users to build, deploy, and manage machine learning models with minimal manual intervention. Designed for both data scientists and business analysts, DataRobot offers a variety of automated features such as model selection, data preprocessing, and hyperparameter tuning. It is particularly useful for those who may not have deep expertise in machine learning but need to leverage the power of AI.
@@ 17.1 Introduction to DataDog
DataDog [92] is a cloud-based monitoring and analytics platform that provides real-time insights into IT infrastructure, applications, and cloud services. It is widely used in industries where monitoring the health of systems, applications, and services is critical to ensuring smooth operations.
@@ 17.3.1 Step 1: Install the DataDog Agent
DD_AGENT_MAJOR_VERSION=7 DD_API_KEY=<YOUR_API_KEY> bash -c "\$(curl -L https://s3.amazonaws.com/dd-agent/scripts/install_script.sh)"
@@ 18.1 Basic Concepts of Artificial Neural Networks
Artificial Neural Networks (ANNs) are a class of algorithms that attempt to mimic the workings of the human brain [93]. The fundamental building blocks of neural networks are neurons (also called nodes or units), which are structured in layers. A neural network typically consists of an input layer, one or more hidden layers, and an output layer [94].
@@
6class SimpleNN(nn.Module):
@@
10 self.fc1 = nn.Linear(2, 3)
11 self.fc2 = nn.Linear(3, 1)
@@ 18.3 Basic Structure of Deep Learning Models
2class DeepNN(nn.Module):
@@
5 self.fc1 = nn.Linear(784, 128) # Input layer (e.g., 28x28 image -> 784)
6 self.fc2 = nn.Linear(128, 64) # First hidden layer
7 self.fc3 = nn.Linear(64, 10) # Output layer (10 classes)
@@ 19.1 Principles of Convolutional Neural Networks
Convolutional Neural Networks (CNNs) [95] are a specialized kind of neural network specifically designed for working with image data, although they can also be applied to other types of structured data. Unlike traditional fully connected networks, CNNs exploit the spatial structure of the data, which is especially useful when dealing with images. In this section, we will walk through the core operations of a CNN: convolution, activation functions, pooling, and fully connected layers.

### v3

@@ 7.2.1 Information Gain
Information gain is a key concept in decision tree learning, where it is used to select the attribute that best separates the data into distinct classes [quinlan1986induction]. It is based on the concept of entropy from information theory, which measures the level of disorder or impurity in a dataset. A split that results in the greatest reduction in entropy is considered the best, as it leads to a more organized and homogenous distribution of classes [mitchell1997machine].
@@
The term entropy has its roots in thermodynamics, where it was used to describe the amount of disorder or randomness in a physical system. The concept was later adapted by Claude Shannon in the 1940s to lay the foundation for information theory, which deals with the transmission, compression, and processing of data [shannon1948mathematical]. Shannon defined entropy as a measure of uncertainty or impurity in a system of information.
In the context of machine learning, entropy quantifies the uncertainty in predicting the class label of an instance in a dataset. If a dataset is perfectly homogeneous (i.e., all instances belong to the same class), the entropy is zero, indicating no uncertainty. On the other hand, if the dataset is evenly split between two or more classes, the entropy is at its maximum, indicating a high degree of uncertainty in classification [tan2006introduction].
@@
[…]
@@
In Shannon’s framework, the goal was to quantify the amount of uncertainty in a message. If you are transmitting a message and the outcome is highly predictable, then little information is gained from receiving it. However, if the outcome is highly uncertain, then the message carries more information. Shannon defined entropy mathematically to quantify this uncertainty [shannon1948mathematical]. In this way, the more uncertain or unpredictable a message is, the more ”information” it contains [mackay2003information].
@@ 7.2.2 Gini Index
The Gini index is a measure used in decision trees, particularly for classification tasks, to evaluate the quality of a split. It assesses the degree of impurity or homogeneity of a node, indicating how well the split separates the data into distinct classes [breiman1984classification]. The Gini index is calculated as:
@@
  • AA is the area between the line of equality and the Lorenz curve.
@@ 7.3 Working Principles of Random Forests
Random Forests are an ensemble learning method that improves decision trees by constructing multiple trees during training and outputting the average prediction of all trees for regression, or the majority vote for classification [breiman2001random]. By using multiple trees, the random forest reduces the risk of overfitting and increases the model’s generalization ability [liaw2002classification].
@@ 8.1 Overview of Boosting Algorithms
Boosting is a powerful ensemble learning technique that combines the predictions of several weak learners to create a strong model [schapire1999brief]. Unlike bagging (such as random forests), where individual models are trained independently, boosting builds models sequentially. Each subsequent model attempts to correct the errors made by the previous ones [freund1997decision]. The goal is to reduce bias, making boosting an effective method for improving model accuracy, especially with complex datasets [natekin2013gradient].
Boosting works by assigning higher weights to misclassified examples, forcing subsequent learners to focus more on these difficult cases. Boosting algorithms have been widely used in competitions like Kaggle due to their high performance, and they are known for being robust to overfitting when tuned correctly [chen2016xgboost].
@@ 8.2 XGBoost
XGBoost stands for eXtreme Gradient Boosting. It is one of the most efficient and scalable implementations of gradient boosting algorithms [chen2016xgboost]. XGBoost has become highly popular due to its speed, performance, and the extensive control it gives over the boosting process.
@@ 8.3 LightGBM
LightGBM (Light Gradient Boosting Machine) is another boosting framework that is optimized for efficiency and speed. It was developed to handle large datasets quickly with lower memory usage [ke2017lightgbm]. LightGBM is known for its ability to handle categorical features natively and for its leaf-wise tree growth strategy, which often results in better accuracy.
@@ 8.4 CatBoost
CatBoost [prokhorenkova2019catboostunbiasedboostingcategorical] is a boosting algorithm developed by Yandex, and it is specifically designed to handle categorical data efficiently. Unlike XGBoost and LightGBM, CatBoost doesn’t require one-hot encoding for categorical features, making it highly effective when working with categorical data.
@@ 9.1.2 Examples of Sparse Models
  • Lasso Regression: A type of linear regression that uses an L1 penalty to shrink some coefficients to zero, effectively removing those features [tibshirani1996regression].
  • Elastic Net: Combines L1 and L2 regularization, creating a balance between ridge regression and lasso regression [zou2005regularization].
@@ 9.2 Principles of Group Lasso
Group Lasso [yuan2006model] is a regularization technique that extends the concept of Lasso by considering groups of features instead of individual ones. It applies L1/L2 norms to groups of variables, leading to entire groups being selected or removed from the model. This is especially useful when the features naturally form groups, such as polynomial features or features derived from categorical variables.
@@ 10.1 Concept of RiskSLIM
RiskSLIM (Risk-Supersparse Linear Integer Models) is a machine learning model designed for tasks that require risk minimization, particularly in contexts where interpretability is critical [ustun2017risk]. Unlike many complex machine learning models that focus on maximizing prediction accuracy at the expense of transparency, RiskSLIM emphasizes both simplicity and predictive performance. It produces highly interpretable scoring systems, often represented as linear models with integer-valued coefficients. These models are particularly useful in domains like healthcare, finance, or law, where decision-making is sensitive, and understanding the model’s reasoning is as important as the accuracy itself [rudin2019stop].
The core idea of RiskSLIM is to balance risk minimization with the constraints of producing sparse, simple models. This balance is achieved by optimizing a loss function while enforcing constraints on the coefficients of the model. The coefficients are typically constrained to be small integers, making the final model easy to interpret and apply, even by non-experts [ustun2019optimizing].
RiskSLIM is ideal when you need to make decisions based on a risk score. For instance, in a medical setting, doctors might use a RiskSLIM model to assess the likelihood of a patient developing a condition based on various health indicators. The output would be a simple, interpretable risk score that directly correlates with the probability of the condition [caruana2015intelligible].
@@ 11.1 Basics of GridSearchCV
Grid search is one of the most popular techniques for hyperparameter tuning in machine learning [bergstra2011algorithms]. It involves exhaustive searching through a manually specified subset of the hyperparameter space of a learning algorithm. When using GridSearchCV, we systematically work through different combinations of parameter values, cross-validating as we go to determine which combination gives the best performance [pedregosa2011scikit].
This method is particularly useful when the search space is small and manageable, as it guarantees that all possible combinations are tested. However, for very large search spaces, more efficient techniques like random search [bergstra2012random] or Bayesian optimization [snoek2012practical] may be preferred.
@@ 12.1 Concept of AutoML
Automated Machine Learning, commonly referred to as AutoML, is the process of automating the end-to-end process of applying machine learning (ML) to real-world problems. Traditionally, applying machine learning techniques [feng2024deeplearningmachinelearning] to a problem requires expertise in data science, programming, and the ability to design machine learning models. However, AutoML aims to simplify this process, making it more accessible to non-experts and reducing the time required to build ML models [he2021automl].
@@
14class SimpleNet($n$.Module):
@@
17 self.fc1 = $n$.Linear(10, 50)
18 self.fc2 = $n$.Linear(50, 2)
@@ 12.2.1 Milestones in the Development of AutoML
  • Early 2010s: The first generation of AutoML tools emerged. Tools such as Auto-WEKA (2013) and Auto-sklearn (2015) [feurer2015efficient] provided early platforms for automating the machine learning pipeline. These focused on automating the model selection and hyperparameter tuning processes.
  • 2018 - Neural Architecture Search (NAS): AutoML evolved beyond classical ML into deep learning with the introduction of NAS. Google introduced NASNet, which automated the process of searching for the best neural network architecture [zoph2018learning]. This was a significant breakthrough in designing efficient neural networks.
  • 2020 - Democratization of AutoML: Many cloud providers began offering AutoML as a service. Tools such as Google AutoML, Azure AutoML, and Amazon SageMaker Autopilot simplified the process of applying AutoML on scalable cloud infrastructure [li2020cloud].
  • Present: Recent advancements in AutoML focus on efficiency, model explainability, and lowering the compute cost of model selection. Research has also been growing around topics like Few-Shot Learning, Zero-Shot Learning, and Hyperparameter Optimization [yao2018taking].
@@ 13.1 Introduction to TPOT
TPOT (Tree-based Pipeline Optimization Tool) is an open-source library designed to automate the process of machine learning, with a particular emphasis on automating feature engineering, model selection, and hyperparameter tuning [olson2016tpot]. It utilizes genetic programming to search for the best possible machine learning pipeline by evaluating different models and combinations of preprocessing steps, features, and hyperparameters.
Machine learning can often be challenging for beginners, as it requires not only a good understanding of the data but also selecting the right models and hyperparameters. TPOT simplifies this process by automatically searching through a range of models and pipeline combinations, thus allowing users to focus more on understanding their data and less on the technical details of model selection and tuning [lam2017automation].
TPOT is built on top of the popular machine learning library scikit-learn [pedregosa2011scikit], and it integrates seamlessly with PyTorch for deep learning tasks. TPOT handles the process of pipeline optimization by evolving the best possible model pipeline over a series of iterations. Each iteration involves generating new pipelines, evaluating their performance, and selecting the best candidates for further optimization. This process continues until a pre-defined number of generations is reached or the optimal solution is found [olson2016evo].
@@ 14.1 Introduction to AutoGluon
AutoGluon [agtabular] is an open-source toolkit designed to simplify machine learning model development. It automates the process of building, tuning, and deploying models by handling most of the complex steps automatically. This makes it an excellent choice for beginners and professionals alike, as it reduces the need for manual tuning of hyperparameters or selecting the best model. Instead, AutoGluon focuses on performance and ease of use, allowing developers to achieve state-of-the-art results without extensive knowledge of machine learning.
@@ 15.1 H2O AutoML
H2O AutoML [H2OAutoML20] is an open-source machine learning platform that automates the entire model training process. It’s known for providing an easy-to-use interface for training and tuning models without needing in-depth expertise in machine learning. With H2O AutoML, you can automatically train a large variety of models, including deep learning, tree-based models, and ensembles. The tool is built to handle both regression and classification problems.
@@ 15.2 MLBox
MLBox is a Python-based AutoML library designed to automate the end-to-end process of machine learning, from preprocessing to model training and optimization [mlbox]. It emphasizes simplicity and automation, making it a great choice for beginners. It also provides powerful preprocessing capabilities, especially for handling missing data and unbalanced datasets [ramirez2018mlbox].
@@ 15.3 Auto-sklearn
  • Built on scikit-learn: Uses the familiar scikit-learn API [pedregosa2011scikit], making it easy to integrate into existing Python workflows.
@@ 15.4 FLAML
FLAML (Fast and Lightweight AutoML) [wang2021flamlfastlightweightautoml] is a lightweight AutoML library developed by Microsoft, focusing on fast and efficient hyperparameter optimization. FLAML is designed to be computationally efficient, making it suitable for users who want to train models quickly without consuming significant computational resources. It is a great tool for handling both classification and regression tasks with a focus on performance and speed.
@@ 16.1 Introduction to DataRobot
DataRobot [datarobot] is a powerful cloud-based AutoML (Automated Machine Learning) platform that allows users to build, deploy, and manage machine learning models with minimal manual intervention. Designed for both data scientists and business analysts, DataRobot offers a variety of automated features such as model selection, data preprocessing, and hyperparameter tuning. It is particularly useful for those who may not have deep expertise in machine learning but need to leverage the power of AI.
@@ 17.1 Introduction to DataDog
DataDog [datadog2021platform] is a cloud-based monitoring and analytics platform that provides real-time insights into IT infrastructure, applications, and cloud services. It is widely used in industries where monitoring the health of systems, applications, and services is critical to ensuring smooth operations.
@@ 17.3.1 Step 1: Install the DataDog Agent
DD_AGENT_MAJOR_VERSION=7 DD_API_KEY=<YOUR_API_KEY> bash -c "\$(curl -L https://s3.amazonaws.com/$d$-agent/scripts/install_script.sh)"
@@ 18.1 Basic Concepts of Artificial Neural Networks
Artificial Neural Networks (ANNs) are a class of algorithms that attempt to mimic the workings of the human brain [lecun2015deep]. The fundamental building blocks of neural networks are neurons (also called nodes or units), which are structured in layers. A neural network typically consists of an input layer, one or more hidden layers, and an output layer [goodfellow2016deep].
@@
6class SimpleNN($n$.Module):
@@
10 self.fc1 = $n$.Linear(2, 3)
11 self.fc2 = $n$.Linear(3, 1)
@@ 18.3 Basic Structure of Deep Learning Models
2class DeepNN($n$.Module):
@@
5 self.fc1 = $n$.Linear(784, 128) # Input layer (e.g., 28x28 image -> 784)
6 self.fc2 = $n$.Linear(128, 64) # First hidden layer
7 self.fc3 = $n$.Linear(64, 10) # Output layer (10 classes)
@@ 19.1 Principles of Convolutional Neural Networks
Convolutional Neural Networks (CNNs) [oshea2015introductionconvolutionalneuralnetworks] are a specialized kind of neural network specifically designed for working with image data, although they can also be applied to other types of structured data. Unlike traditional fully connected networks, CNNs exploit the spatial structure of the data, which is especially useful when dealing with images. In this section, we will walk through the core operations of a CNN: convolution, activation functions, pooling, and fully connected layers.